# Approximating the coefficients of the Bessel functions

### v4

@@ Abstract.
For the type A, BC, and D root systems, we determine equivalent conditions between the coefficients of an exponential holomorphic function and the asymptotic values taken by the Dunkl bilinear form when one of its entries is the function. We establish these conditions over the $|\theta N|\rightarrow\infty$ regime for the type A and D root systems and over the $|\theta_{0}N|\rightarrow\infty,\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ regime for the type BC root system. We also generalize existing equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type A root system and over the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ regime for the type BC root system and prove new equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type D root system. Furthermore, we determine the asymptotics of the coefficients of the Bessel functions over the regimes that we have mentioned.
@@ 1. Introduction
For positive integers $N\geq 2$, we study the asymptotics of the Bessel functions associated to the irreducible root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. We denote the Bessel function associated to a finite root system $\mathcal{R}$ and multiplicity function $\theta$ as $J^{\mathcal{R}(\theta)}_{a}(x)$, where $a,x\in\mathbb{C}^{N}$, provided that the function exists. It is an eigenfunction of the Dunkl operators associated to $\mathcal{R}$ and $\theta$, which are introduced in [dunkloperators]. Furthermore, the function is symmetric with respect to the reflection group generated by $\mathcal{R}$ in the variables $a$ and $x$. For the definitions of root systems, multiplicity functions, and the Bessel functions, see Subsection˜2.1.
For certain choices of $\theta$, the paper [dunklbessel] shows that $J^{\mathcal{R}(\theta)}_{a}(x)$ exists, is unique, and is holomorphic over $a,x\in\mathbb{C}^{N}$, see Theorem˜2.8. The paper also discusses a nonsymmetric eigenfunction, see Theorem˜2.5; the formula relating the two eigenfunctions is given in Theorem˜2.8.
In this paper, we analyze the coefficients of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is fixed and $\mathcal{R}$ is one of $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. The multiplicity function $\theta$ varies with $N$ and should be viewed as a function of $N$, although we denote it as $\theta$ rather than $\theta(N)$ for brevity. We outline the notation we use to record the root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$ and the corresponding multiplicity functions. Note that $e_{i}\triangleq[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}$ for $i\in[N]$.
@@
for all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are homogeneous, have the same degree, and symmetric with respect to the reflection group generated by $\mathcal{R}$; for the explanation of why this is the case, see Lemma˜4.19. We compute the asymptotics of these values in Theorem˜1.1.
@@ 1.1. Main result
For the notation regarding partitions and noncrossing partitions in the following result, see Subsections˜2.2 and 2.3, respectively. We briefly mention that for $N\geq 1$, $\Gamma_{N}$ consists of partitions with all parts of size at most $N$, $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}$, and $\Gamma_{\text{even}}$ and $\Gamma_{\text{even};\,N}$ are the elements of $\Gamma$ and $\Gamma_{N}$, respectively, with all parts of even size; for $\lambda\in\Gamma$, $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$. Additionally, $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$. Furthermore, for $k\geq 1$, $NC(k)$ denotes the set of of noncrossing partitions of $[k]$ and $NC^{\text{even}}(k)$ denotes the set of noncrossing partitions of $[k]$ with all even block sizes.
@@
Then, (c) implies (d), and if $c\not=-1$, then (d) implies (c).
@@
See Corollaries˜6.11, 7.5 and 8.22 for the proofs of (A), (B), and (C), respectively. ∎
First, we note that we actually prove the generalizations where the $N\rightarrow\infty$ limits of $\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$, $\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}$, and $\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$ can be nonzero when $\ell(\lambda)>1$, see Theorems˜6.10, 7.4 and 8.21. We present this corollary because it is more applicable to the setting where the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are set as $J_{a(N)}^{A^{N-1}(\theta)}$, $J_{a(N)}^{BC^{N}(\theta_{0},\theta_{1})}$, and $J_{a(N)}^{D^{N}(\theta)}$ in some sufficiently small neighborhood of the origin where the Bessel functions are nonzero so that their logarithms are holomorphic over the neighborhood; $a(N)\in\mathbb{C}^{N}$ varies with $N$.
@@
Then, we can deduce the convergence of the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ in terms of moments. Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin. We prove a generalized version of this result in Corollary˜10.17.
@@
Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]; see Lemma˜1.2 for a more direct connection with the question that is posed. In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$. The main contribution of this paper is the set of equivalence relations in the theorem where some conditions are based on scaling the coefficients of the logarithms of the input functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ by varying powers of $\theta N$. The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed is mentioned in the appendix of [matrix]. However, in [limits_prob_measures], the coefficients are only scaled by $(\theta N)^{-1}$ and it is similarly the case that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
@@
As we mention in Section˜6, a method to prove part (A) of Theorem˜1.1 is to prove Theorem˜6.1 using the results of [limits_prob_measures]. However, this method is not applicable to parts (B) and (C), so we develop a new approach which can be used in these settings in Subsection˜6.1. This approach is applicable to the $\theta N\rightarrow c\in\mathbb{C}$ regime as well.
Interestingly, in part (C) of Theorem˜1.1, we also consider the asymptotics of coefficients of terms with degree increasing to infinity. See Subsection˜1.6, Section˜8, and Section˜11 for more discussion regarding this direction.
We can also restate Theorem˜1.1 in terms of the Taylor series expansions over the standard basis $M_{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ and the coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$. See Subsection˜2.2 for the definitions of these notions.
Lemma 1.2.
Repeat the statement of Theorem˜1.1. Suppose
$F_{N}^{A}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m^{A}_{\epsilon}(N)M_{\epsilon})\right)\in\mathbb{C}[[x_{1},\ldots,x_{N}]],$
$F_{N}^{BC}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{BC}_{\epsilon}(N)M_{\epsilon}\right)\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]],$
$F_{N}^{D}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{D}_{\epsilon}(N)M_{\epsilon}\right)+e\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}l^{D}_{\epsilon}(N)M_{\epsilon}\right)$
$\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]].$
Consider the following statements.
  1. (a’)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (c’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (e’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  4. (f’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
Then, (a’), (c’), (e’), and (f’) are equivalent to (a), (c), (e), and (f), respectively.
Remark 1.3.
The partial derivative $\partial^{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ denotes $\prod_{i=1}^{\ell(\epsilon)}\partial_{i}^{\epsilon_{i}}$. We can clearly replace it with $\prod_{i=1}^{\ell(\epsilon)}\partial_{\pi(i)}^{\epsilon_{i}}$ for any injection $\pi:[\ell(\epsilon)]\rightarrow[N]$ since the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are symmetric. Furthermore, in (f”), $(1,\ldots,1)$ consists of $N$ copies of $1$ and $(\epsilon+(1,\ldots,1))!=\prod_{i=1}^{\ell(\epsilon)}(\epsilon_{i}+1)$.
Proof of Lemma˜1.2.
The equivalences are straightforward after using the definition of the coefficients $c_{\epsilon\lambda}$. ∎
1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
Observe that in Theorem˜1.1, we consider the following regimes:
@@ 1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
Note that [matrix] considers regime 3 for the root system $A^{N-1}$ while [rectangularmatrix] considers regime 4. The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well. Theorems that we prove over regimes 3 and 4 include Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix]. We also prove Theorem˜8.26 for regime 3 and the type D root system.
1.3. Related works
Some related works that we have already mentioned are [matrix, limits_prob_measures, rectangularmatrix]. These works are based on analyzing the Bessel generating functions of exponentially decaying measures to compute the asymptotic moments of the measures. While we focus on a broader setting, we discuss the applications of this paper’s results to exponentially decaying measures in Subsection˜10.5. Another work that is based on analyzing Bessel generating functions to compute asymptotic moments is [gaussianfluctuation], which considers when $\theta=1$. Furthermore, the papers [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov] consider similar results by analyzing the coefficients of Jack and Schur generating functions.
The paper [limits_prob_measures] introduces a framework for analyzing the asymptotic moments when the coefficients of terms of the Taylor series of the type A Bessel generating functions with two or more variables can have nonzero $N\rightarrow\infty$ limits after scaling by $(\theta N)^{-1}$. In fact, we are able to apply the results of the paper to obtain a straightforward proof of part (A) of Theorem˜1.1, as we mentioned earlier. In addition, we present a new proof method that is applicable to other settings, such as the $BC^{N}$ and $D^{N}$ root systems and regimes 3 and 4 that we described in Subsection˜1.2.
1.4. Weak convergence to the free convolution
Note that the definitions of $\theta(\mathcal{R})$ and $H(\mathcal{R})$ are included in Subsection˜2.1.
First, we discuss an integral representation of the nonsymmetric eigenfunction for the Dunkl operators. This integral representation assumes that the multiplicity function is nonnegative.

### v5

@@ Abstract.
We determine equivalent conditions between the asymptotic coefficients of the Bessel generating functions of a sequence of probability measures and the asymptotic expected values of power sums when their inputs are sampled from these measures. We establish these conditions over the $|\theta N|\rightarrow\infty$ regime for the type A and D root systems and over the $|\theta_{0}N|\rightarrow\infty,\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ regime for the type BC root system. We also establish equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type A and D root systems and over the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ regime for the type BC root system that generalize existing results. Furthermore, we determine the asymptotics of the coefficients of the Bessel functions over the regimes that we have mentioned.
@@ 1. Introduction
For positive integers $N\geq 2$, we study the asymptotics of the Bessel functions associated to the irreducible root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. We denote the Bessel function associated to a finite root system $\mathcal{R}$ and multiplicity function $\theta$ by $J^{\mathcal{R}(\theta)}_{a}(x)$, where $a,x\in\mathbb{C}^{N}$, provided that the function exists. It is an eigenfunction of the Dunkl operators associated to $\mathcal{R}$ and $\theta$, which are introduced in [dunkloperators]. Furthermore, the function is symmetric with respect to the reflection group generated by $\mathcal{R}$ in the variables $a$ and $x$.
For certain choices of $\theta$, the paper [dunklbessel] shows that $J^{\mathcal{R}(\theta)}_{a}(x)$ exists, is unique, and is holomorphic over $a,x\in\mathbb{C}^{N}$, see Theorem˜2.8. The paper also discusses a nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$, see Theorem˜2.5; we have that $J^{\mathcal{R}(\theta)}_{a}(x)$ is the average of $E_{ga}^{\mathcal{R}(\theta)}(x)$ when $g$ is selected uniformly at random from the finite reflection group generated by $\mathcal{R}$.
In this paper, we analyze the coefficients of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is fixed and $\mathcal{R}$ is one of $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. The multiplicity function $\theta$ varies with $N$ and should be viewed as a function of $N$, although we denote it as $\theta$ rather than $\theta(N)$ for brevity.
for all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are homogeneous, have the same degree, and symmetric with respect to the reflection group generated by $\mathcal{R}$; for an explanation of why this is the case, see Lemma˜4.19. We compute the asymptotics of these values in Theorem˜1.2.
1.1. Notation
We introduce some notation required for stating the main result. For additional definitions and notation, see Section˜2.
For $N\geq 1$, $\Gamma_{N}$ consists of partitions with all parts of size at most $N$, $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}$, and $\Gamma_{\text{even}}$ and $\Gamma_{\text{even};\,N}$ are the elements of $\Gamma$ and $\Gamma_{N}$, respectively, with all parts of even size; for $\lambda\in\Gamma$, $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$. Additionally, $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$. Furthermore, for $k\geq 1$, $NC(k)$ denotes the set of noncrossing partitions of $[k]$ and $NC^{\text{even}}(2k)$ denotes the set of noncrossing partitions of $[2k]$ with all even block sizes.
We outline the notation we use to denote the root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$ and the corresponding multiplicity functions. Note that $e_{i}\triangleq[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}$ for $i\in[N]$.
@@ 1.1. Notation
1.2. Bessel generating functions and related works
In addition to considering the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ for a single value of $a$, we can consider the average of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is randomly sampled. The resulting function is referred to as the Bessel generating function and has been studied previously in [gaussianfluctuation, matrix, airy_beta, limits_prob_measures, rectangularmatrix]. This notion is also related to the Dunkl transform introduced in [dunklbound]. We discuss results related to Bessel generating functions in Subsection˜10.5.
Definition 1.1.
Suppose $\mu$ is a Borel probability measure over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
$G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\mathbb{E}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)],$
provided that the expected value converges.
For $k\geq 1$, the type A Bessel function satisfies
$\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}J_{a}^{A^{N-1}(\theta)}(x)=\sum_{i=1}^{N}a_{i}^{k}J_{a}^{A^{N-1}(\theta)}(x).$
Therefore, after assuming regularity conditions on $\mu$, we have that
$1]\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}G_{\mu}^{A^{N-1}(\theta)}(x)=\mathbb{E}_{a\sim\mu}\left[\sum_{i=1}^{N}a_{i}^{k}\right],$
see Lemma˜10.16 for more details. As $N\rightarrow\infty$, in order to analyze the moments of $\mu$, it suffices to analyze $1]\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}G_{\mu}^{\mathcal{R}(\theta)}(x)$. For the type A root system, the paper [gaussianfluctuation] addresses this question for $\theta=1$ and [matrix] does so for the $\theta N\rightarrow c\geq 0$ regime. For the type BC root system, [rectangularmatrix] does so for the $\theta_{0}N\rightarrow c_{0}\geq 0$ and $\theta_{1}\rightarrow c_{1}\geq 0$ regime. We address this question for both regimes and the type A, BC, and D root systems.
The paper [gaussianfluctuation] defines the Bessel generating function to be the average of $\frac{J_{a}^{\mathcal{R}(\theta)}(x)}{J_{a}^{\mathcal{R}(\theta)}(\rho)}$ when $a$ is sampled from a probability ensemble; we consider when $\rho=0$. In order to determine the moments of the probability ensemble, the paper computes the value of the Bessel generating function after applying Dunkl operators when $x=\rho$ and as $N\rightarrow\infty$, $\rho$ converges to a fixed distribution after rescaling.
For the $|\theta N|\rightarrow\infty$ regime, the paper [limits_prob_measures] introduces a framework for analyzing the asymptotic moments when the coefficients of the logarithm of the type A Bessel generating functions with two or more variables can have nonzero $N\rightarrow\infty$ limits after scaling by $(\theta N)^{-1}$. The paper [domino_tiling] studies Schur generating functions and considers the same setting.
The converse statement is that we can determine the asymptotics of the coefficients of $G_{\mu}^{\mathcal{R}(\theta)}$ as $N\rightarrow\infty$ from the values of $p_{\nu}(\mu)$ for $\nu\in\Gamma$. The paper [matrix] establishes this statement for the type A root system and the $\theta N\rightarrow c\geq 0$ regime and asks whether it is true for the $|\theta N|\rightarrow\infty$ regime. We answer this question and establish the converse statement for both regimes and the type A, BC, and D root systems.
Furthermore, the papers [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov, domino_tiling] similarly analyze the coefficients of Jack and Schur generating functions to establish a law of large numbers. The papers [unitary, llnclt, jackgenfunc, rectangularmatrix] also prove converse statements and [fluctuations, unitary, llnclt, gaussianfluctuation, domino_tiling] consider the central limit theorem in addition to the law of large numbers.
@@ 1.3. Main result
Then, (c) implies (d) and if $c\not=-1$, then (d) implies (c).
@@
See Corollaries˜6.11, 7.5 and 8.10 for the proofs of (A), (B), and (C), respectively. ∎
First, we note that we actually prove the generalizations where the $N\rightarrow\infty$ limits of $\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$, $\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}$, and $\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$ can be nonzero when $\ell(\lambda)>1$, see Theorems˜6.10, 7.4 and 8.9. We present this corollary because it is more applicable to the setting where the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are set as $J_{a(N)}^{A^{N-1}(\theta)}$, $J_{a(N)}^{BC^{N}(\theta_{0},\theta_{1})}$, and $J_{a(N)}^{D^{N}(\theta)}$ in some sufficiently small neighborhood of the origin where the Bessel functions are nonzero so that their logarithms are holomorphic over the neighborhood; $a(N)\in\mathbb{C}^{N}$ varies with $N$.
@@
Then, we can deduce the convergence of the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 2}$ in terms of moments. Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 2}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin. We prove a generalized version of this result in Corollary˜10.18.
@@
As we mentioned previously, Theorem˜1.2 resolves the question posed in the appendix of the paper [matrix]. The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed that we use is mentioned in the paper. In particular, we generalize the results of [limits_prob_measures], where the coefficients are only scaled by $(\theta N)^{-1}$.
@@
As we mention in Section˜6, a method to prove part (A) of Theorem˜1.2 is to prove Theorem˜6.1 using the results of [limits_prob_measures]. However, this method is not applicable to parts (B) and (C), so we develop a new approach which can be used in these settings in Subsection˜6.1. This approach is applicable to the $\theta N\rightarrow c\in\mathbb{C}$ regime as well.
1.4. The high temperature regime
Observe that in Theorem˜1.2, we consider the following regimes:
@@ 1.4. The high temperature regime
The approach that we use to prove Theorem˜1.2 over regimes 1 and 2 is adaptable to proving similar results over regimes 3 and 4 as well. Theorems that we prove over regimes 3 and 4 include Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix]. We also prove Theorem˜8.11 for regime 3 and the type D root system.
1.5. Proof method
For the proof of Theorem˜1.2 for the type A root system, the main idea is that when $|\theta N|\rightarrow\infty$ and $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is applied to $p_{\lambda}$ for some $\lambda\in\Gamma$, only certain sequences of operators contribute to the leading order term. Initialize $S$ to be $\ell(\lambda)]$. In order to contribute to the leading order term, we must have that the first operator of $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is $\partial_{i}$ applied to $p_{(\lambda_{j})}$ for some $j\in S=[\ell(\lambda)]$, since $p_{\lambda}$ is symmetric. Afterwards, we remove $j$ from $S$ and replace $p_{\lambda}$ with $\lambda_{j}x_{i}^{\lambda_{j}-1}\frac{p_{\lambda}}{p_{(\lambda_{j})}}$. For subsequent operators, we either diminish the degree of $x_{i}$ by one and multiply by $\theta N$ or apply $\partial_{i}$ to $p_{(\lambda_{j})}$ for some $j\in S$ and remove $j$ from $S$. Afterwards, we can deduce the result by constructing a bijection with noncrossing partitions.
Let $d_{i}$ denote the operator which diminishes the degree of $x_{i}$ by one. Then, to contribute to the leading order term, $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is equivalent to $(\partial_{i}+\theta Nd_{i})^{k-1}\partial_{i}$, where $\partial_{i}$ is only applied to $p_{(\lambda_{j})}$ for $j\in S$.
[…]
In the $\theta N\rightarrow c\in\mathbb{C}$ regime, we have that $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is equivalent to $(\partial_{i}+cd_{i})^{k-1}\partial_{i}$. In contrast, we do not impose the constraint that $\partial_{i}$ is only applied to $p_{(\lambda_{j})}$ for $j\in S$. Similarly, in the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$ and $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime, $\mathcal{D}_{i}(BC^{N}(\theta_{0},\theta_{1}))^{k}$ is equivalent to $\left(\partial_{i}+2c_{0}d_{i}+c_{1}\frac{1-\tau_{i}}{x_{i}}\right)^{k-1}\partial_{i}$. Afterwards, we similarly construct a bijection with noncrossing partitions.
1.6. Probabilistic applications
As examples of applications, we prove that an observation of the DBM converges to a fixed measure as $\theta N\rightarrow\infty$ and determine the measure; it is straightforward to prove an analogous result in the $\theta N\rightarrow c\geq 0$ regime. We can similarly establish the law of large numbers for the $\beta$-Hermite ensemble as $\theta N\rightarrow\infty$, see [limits_prob_measures]; this ensemble is equivalent to the DBM started at zero. For the $\beta$-Laguerre ensemble, we also discuss the law of large numbers as $\theta N\rightarrow\infty$ and reprove a result of [betalaguerre].
Although we only consider the type A DBM in this paper, we can prove analogous results for the type BC and type D DBMs using the same method. For the definitions of these Brownian motions, see [demni_sde].
The paper [airy_beta] also uses Dunkl operators to analyze the type A DBM. The paper shows that multiple observations of the largest particles of the DBM converge to the Airyβ line ensemble.
1.7. Integral representations of the Bessel functions
First, we introduce the intertwining measure, which is a representation of the Bessel function as an integral of exponential functions. For a finite root system $\mathcal{R}$, $H(\mathcal{R})$ denotes the finite reflection group generated by $r_{\alpha}$ for $\alpha\in\mathcal{R}$ and for $a\in\mathbb{R}^{N}$, $H(\mathcal{R})a\triangleq\{ha:h\in H(\mathcal{R})\}$. Moreover, $\theta(\mathcal{R})$ is the set of multiplicity functions over $\mathcal{R}$.